# On functoriality and the tensor product property in noncommutative tensor-triangular geometry

### v1

@@ 6. Intermezzo: Quasi-support data
While we are on the topic of finding the universal support datum for a given set of support axioms, we may as well address quasi support data, as introduced in [NVY22a, Section 7.1].
@@
First, we must verify that $\operatorname{Sp}^{r}(\mathscr{K})$ is a quasi support datum. Verifications of (a)-(d) are standard. For (e), we have for any $\mathscr{I}\in\operatorname{Sp}^{r}(\mathscr{K})$ that $x\in\mathscr{K}$ implies $x\otimes y\in\mathscr{K}$, hence $\mathscr{I}\in\sup(x\otimes y)\implies\mathscr{I}\in\sup(x)$, thus $(\operatorname{Sp}(\mathscr{K}),\sup)$ is a quasi support datum. Now, given any quasi support datum $(Y,\sigma)$, we must check that $f$ is well-defined, continuous, unique, and it satisfies $\sigma(a)=f^{-1}(\sup(a)$. This is a familiar drill by now: uniqueness follows by the same argument as in 4.4. To check $f$ is well-defined, it suffices to verify that $f(y)$ is a thick right $\otimes$-ideal. Let $a\in f(y)$, then $y\not\in\sigma(a)$. Condition (e) implies $y\not\in\sigma(a\otimes b)$ for any $b\in\mathscr{K}$, therefore $a\otimes b\in f(y)$ for any $b\in\mathscr{B}$, as desired. By definition, we have $f(y)\in\sup(a)$ if and only if $a\not\in f(y)$ if and only if $y\in\sigma(a)$. Therefore, $f^{-1}(\sup(a))=\sigma(a)$, as desired, and it follows that $f$ is continuous. ∎
@@
Not to despair, for [NVY22a, Theorem 7.3.1] still provides a way of classifying thick right $\otimes$-ideals (i.e. determining $\operatorname{Sp}(\mathscr{K})$) via support methods, although per remark [NVY22a, Remark 7.3.2], this technique only works when one’s quasi-support datum is in fact a noncommutative support datum. However, the compression of information into the prime right $\otimes$-ideals does not happen for the universal support. Quoting Balmer and Ocal: “This is not because of whimsical choices in the construction. Being the solution to a universal problem, it is what it is.”
@@ 7. The complete prime spectra of crossed product categories
Finally, for continuity, if we denote the bijection $\Phi\colon\operatorname{Spc}^{cp}(\mathscr{K}\rtimes G)\to G\operatorname{-}\operatorname{Spc}^{cp}(\mathscr{K}),$ then it is straightforward to verify that $\Phi^{-1}(\operatorname{supp}^{cp}(x))=\operatorname{supp}^{cp}(x\boxtimes g)$ for every $x\in\mathscr{K}$ and $g\in G$. Since the supports are a set of closed generators of the completely prime spectrum, $\Phi$ is a homeomorphism, as desired. ∎
@@
It is easy to see the above example generalizes to any finite group $G$ and essentially small tensor-triangulated category $\mathscr{K}$ with $G$ acting on $\mathscr{K}^{\prime}:=\mathscr{K}^{\oplus|G|}$ by permuting indices, a categorical “wreath product” if one will.
@@ 8. Duality and duo-ness
We have one general criterion for right duo-ness for well-behaved finite dimensional Hopf algebras. This is used in the computations in [NVY22b, Section 5]; we record it for completeness. Note that by [NVY24], one always has a homomorphism into Proj of the categorical center of the cohomology ring, but it is by no means a homeomorphism in general.
@@
  • For all $M,N\in\operatorname{stab}(A)$, $\operatorname{Ext}^{*}(M,N)$ is a finitely generated $\operatorname{H}^{*}(A,k)$-module;
@@
then $\operatorname{stab}(A)$ is duo.
@@
The finite generation conditions are conjectured to hold for all finite tensor categories by Etingof and Ostrik, see [EO04]. The conditions of Proposition 8.8 have been shown to hold in a myriad of settings, see [FP07] for the case of finite group schemes, and [NVY22a] and [NP23] for various noncommutative settings. The given homeomorphism on Balmer spectra is conjectured to hold for all quasi-triangular Hopf algebras by Nakano, Vashaw, and Yakimov, see [NVY24] (more generally, they predict a modification of this homeomorphism holding for all finite tensor categories). If both conjectures hold, it follows that $\operatorname{stab}(A)$ is duo for any finite dimensional quasi-triangular Hopf algebra $A$. Their conjecture would also provide a positive answer to a question asked by Bergh, Plavnik, and Witherspoon in [NVY24] on whether the cohomological support of a braided finite tensor category with finitely generated cohomology has the tensor product property.
@@ 9. Surjectivity and tensor-nilpotence
There are two key points which prevent us from fully extending [Bal18, Theorem 1.4], a stronger statement that characterizes exactly when $\varphi:\operatorname{Spc}(\mathscr{L})\to\operatorname{Spc}(\mathscr{K})$ is surjective in generality. In one direction (the direction (a) $\implies$ (b) in the proof), it requires that the collection $\{z\in\mathscr{K}\mid f\text{ is $\otimes$-nilpotent on }z\}$ be a thick triangulated $\otimes$-ideal, which as noted in 9.2, may not be a triangulated subcategory. Additionally, in the (b) $\implies$ (a) direction, it is deduced from the fact that $\xi_{x}^{\otimes n}$ is $\otimes$-nilpotent on its cone that $g=\xi_{x}^{\otimes n}\otimes s$ is $\otimes$-nilpotent on its cone as well; this does not hold in our case, even if we replace $\otimes$-nilpotent” with “weakly $\otimes$-nilpotent.”
@@ 10. Finale: the universal functorial support theory
These subcategories are very ad-hoc - usually one only considers strictly-full subcategories of a triangulated category. The point is only that they exist and that their inclusion is an exact $\otimes$-functor. We now give an important reformulation of partial primeness.
@@
Let $\mathscr{P}$ be a full subcategory of $\mathscr{K}$. The following are equivalent: $\mathscr{P}$ is a partial prime $\otimes$-ideal of $\mathscr{K}$ if and only if for every (not necessarily full or thick) braided monoidal triangulated subcategory $\mathscr{C}$ of $\mathscr{K}$, $\mathscr{P}\cap\mathscr{C}$ is a prime thick $\otimes$-ideal of $\mathscr{C}$.
@@
In particular, the partial $\otimes$-ideal consisting of all nilpotent elements of $\mathscr{K}$ is contained in all prime partial $\otimes$-ideals.
@@
Let $\operatorname{Br}(\mathscr{K})$ denote the set of all braided monoidal triangulated subcategories of $\mathscr{K}$. In fact, $\operatorname{Br}(\mathscr{K})$ is a poset under inclusion, and because inclusion of monoidal triangulated subcategories is an exact $\otimes$-functor, $\operatorname{Br}(\mathscr{K})$ is a sub-poset of $\mathbf{Bmon}_{\Delta}$, the category of essentially small braided monoidal triangulated categories.
@@
One may carry out the same process with the phrase “symmetric” replacing “braided,” but with the additional imposed condition that for $x,y$ to “commute,” there must exist an isomorphism $\tau:x\otimes y\cong y\otimes x$ such that $\tau^{2}=\operatorname{id}$.
@@
The most pertinent question that follows is whether there exists a monoidal triangulated category for which $p\operatorname{Spc}(\mathscr{K})=\emptyset$. At present we are not aware of any such categories. One can attempt to recreate a ring $R$ as an ad-hoc monoidal “triangulated” category by setting objects the elements of $R$ and only identity morphisms, with direct sums corresponding to addition and tensor products corresponding to multiplication. The set of prime thick $\otimes$-ideals of this category does not coincide with the prime ideals of $R$, however, due to thickness.
Another question of interest is whether there exists a monoidal triangulated category $\mathscr{K}$ for which $p\operatorname{Spc}(\mathscr{K})=\operatorname{Spc}^{cp}(\mathscr{K})\neq\emptyset$. The first equality obviously holds for any category $\mathscr{K}$ for which $p\operatorname{Spc}(\mathscr{K})=\emptyset$.

### v2

@@ 7. Intermezzo: Quasi-support data
While we are on the topic of finding the universal support datum for a given set of support axioms, we address quasi support data, as introduced in [NVY22a, Section 7.1].
@@
First, we must verify that $\operatorname{Sp}^{r}(\mathscr{K})$ is a quasi support datum. Verifications of (a)-(d) are standard. For (e), we have for any $\mathscr{I}\in\operatorname{Sp}^{r}(\mathscr{K})$ that $x\in\mathscr{K}$ implies $x\otimes y\in\mathscr{K}$, hence $\mathscr{I}\in\sup(x\otimes y)\implies\mathscr{I}\in\sup(x)$, thus $(\operatorname{Sp}(\mathscr{K}),\sup)$ is a quasi support datum. Now, given any quasi support datum $(Y,\sigma)$, we must check that $f$ is well-defined, continuous, unique, and it satisfies $\sigma(a)=f^{-1}(\sup(a)$. This is a familiar drill by now: uniqueness follows by the same argument as in 5.1. To check $f$ is well-defined, it suffices to verify that $f(y)$ is a thick right $\otimes$-ideal. Let $a\in f(y)$, then $y\not\in\sigma(a)$. Condition (e) implies $y\not\in\sigma(a\otimes b)$ for any $b\in\mathscr{K}$, therefore $a\otimes b\in f(y)$ for any $b\in\mathscr{B}$, as desired. By definition, we have $f(y)\in\sup(a)$ if and only if $a\not\in f(y)$ if and only if $y\in\sigma(a)$. Therefore, $f^{-1}(\sup(a))=\sigma(a)$, as desired, and it follows that $f$ is continuous. ∎
@@
Not to despair, for [NVY22a, Theorem 7.3.1] still provides a way of classifying thick right $\otimes$-ideals (i.e. determining $\operatorname{Sp}(\mathscr{K})$) via support methods, although per remark [NVY22a, Remark 7.3.2], this technique only works when one’s quasi-support datum is in fact a noncommutative support datum. Indeed, it seems the lack of well-behavedness of rigidity with one-sided ideals poses a problem for classification. However, the compression of information into the prime right $\otimes$-ideals does not happen for the universal support. Quoting Balmer and Ocal: “This is not because of whimsical choices in the construction. Being the solution to a universal problem, it is what it is.”
@@ 8. The complete prime spectra of crossed product categories
Finally, for continuity, if we denote the bijection $\Phi\colon\operatorname{Spc}^{cp}(\mathscr{K}\rtimes G)\to G\operatorname{-}\operatorname{Spc}^{cp}(\mathscr{K}),$ with inverse $\Psi$, then it is straightforward to verify that
$\Phi^{-1}(\operatorname{supp}^{cp}(x))=\operatorname{supp}^{cp}(x\boxtimes g)$
for every $x\in\mathscr{K}$ and $g\in G$, and similarly,
$\operatorname{supp}^{cp}(x)=\Psi^{-1}(\operatorname{supp}^{cp}(x\boxtimes g)).$
Thus $\Phi$ is a homeomorphism, as desired. ∎
@@
It is easy to see the above example generalizes to any finite group $G$ and essentially small tensor-triangulated category $\mathscr{K}$ with $G$ acting on $\mathscr{K}^{\prime}:=\mathscr{K}^{\oplus|G|}$ by permuting indices, a categorical “wreath product” when $G$ is a symmetric group.
@@ 9. Duality and duo-ness
We note one general criterion for right duo-ness for well-behaved finite dimensional Hopf algebras. This is used in the computations in [NVY22b, Section 5]; we record it for completeness. Note that by [NVY24], one always has a homomorphism into Proj of the categorical center of the cohomology ring, but it is by no means a homeomorphism in general.
@@
  • For all $M\in\operatorname{stab}(A)$, $\operatorname{Ext}^{\scriptscriptstyle\bullet}(M,M)$ is a finitely generated $\operatorname{H}^{\scriptscriptstyle\bullet}(A,k)$-module;
@@
then $\operatorname{stab}(A)$ is right duo.
@@
The finite generation conditions are conjectured to hold for all finite tensor categories by Etingof and Ostrik, see [EO04]. The given homeomorphism on Balmer spectra is conjectured to hold for all quasi-triangular Hopf algebras by Nakano, Vashaw, and Yakimov, see [NVY24] (more generally, they predict a modification of this homeomorphism holding for all finite tensor categories). If both conjectures hold, it follows that $\operatorname{stab}(A)$ is duo for any finite dimensional quasi-triangular Hopf algebra $A$. Their conjecture would also provide a positive answer to a question asked by Bergh, Plavnik, and Witherspoon [NVY24] on whether the cohomological support of a braided finite tensor category with finitely generated cohomology has the tensor product property.
@@ 10. Surjectivity and tensor-nilpotence
There are two key points which prevent us from fully extending [Bal18, Theorem 1.4], a stronger statement that characterizes exactly when $\varphi:\operatorname{Spc}(\mathscr{L})\to\operatorname{Spc}(\mathscr{K})$ is surjective in generality. In one direction (the direction (a) $\implies$ (b) in the proof), it requires that the collection $\{z\in\mathscr{K}\mid f\text{ is $\otimes$-nilpotent on }z\}$ be a thick triangulated $\otimes$-ideal, which as noted in 10.2, may not be a triangulated subcategory. Additionally, in the (b) $\implies$ (a) direction, it is deduced from the fact that $\xi_{x}^{\otimes n}$ is $\otimes$-nilpotent on its cone that $g=\xi_{x}^{\otimes n}\otimes s$ is $\otimes$-nilpotent on its cone as well; this does not hold in our case, even if we replace $\otimes$-nilpotent” with “weakly $\otimes$-nilpotent.”
@@ 11. Finale: the universal functorial support theory
These subcategories are rather ad-hoc - the point is only that they exist and that their inclusion is an exact $\otimes$-functor, since we require that the subcategory shares the same tensor unit. We now give an important reformulation of partial primeness.
@@
Let $\mathscr{P}$ be a full subcategory of $\mathscr{K}$. The following are equivalent: $\mathscr{P}$ is a partial prime $\otimes$-ideal of $\mathscr{K}$ if and only if for every (not necessarily full or thick) braided monoidal triangulated subcategory $\mathscr{C}$ of $\mathscr{K}$ with the same tensor unit, $\mathscr{P}\cap\mathscr{C}$ is a prime thick $\otimes$-ideal of $\mathscr{C}$.
@@
For instance, the partial $\otimes$-ideal consisting of all nilpotent elements of $\mathscr{K}$ is contained in all prime partial $\otimes$-ideals.
@@
Let $\operatorname{Br}(\mathscr{K})$ denote the set of all braided monoidal triangulated subcategories of $\mathscr{K}$ with the same tensor unit. In fact, $\operatorname{Br}(\mathscr{K})$ is a poset under inclusion, and because inclusion of monoidal triangulated subcategories is an exact $\otimes$-functor, $\operatorname{Br}(\mathscr{K})$ is a sub-poset of $\mathbf{Bmon}_{\Delta}$, the category of essentially small braided monoidal triangulated categories.
@@
One may carry out the same process of creating the partial prime spectrum with the phrase “symmetric” replacing “braided,” but with the additional imposed condition that for $x,y$ to “commute,” there must exist an isomorphism $\tau:x\otimes y\cong y\otimes x$ such that $\tau^{2}=\operatorname{id}$.
@@
In fact, any functorial spectral theory $(F,\sigma)$ on essentially small monoidal triangulated categories agreeing with $(\operatorname{Spc},\operatorname{supp})$ on all braided monoidal triangulated categories must also be a commutative support datum on $\mathscr{K}$. Indeed, for any braided monoidal subcategory $\mathscr{C}\subseteq\mathscr{K}$ with the same tensor unit, one has a map on spectra $\operatorname{Spc}(\iota):F(\mathscr{K})\to F(\mathscr{C})\cong\operatorname{Spc}(\mathscr{C})$ satisfying
$\operatorname{Spc}(\iota)^{-1}(\operatorname{supp}(x))=\sigma(\iota(x))=\sigma(x)$
for all $x\in\mathscr{C}$.
The next pertinent question is now: does there exist a monoidal triangulated category $\mathscr{K}$ for which $p\operatorname{Spc}(\mathscr{K})=\emptyset$? At present, we are not aware of such a category, but conjecture such a category exists. We also ask if there exists a non-braided monoidal triangulated category $\mathscr{K}$ for which $p\operatorname{Spc}(\mathscr{K})\cong\operatorname{Spc}^{cp}(\mathscr{K})\neq\emptyset$. Of course, if $p\operatorname{Spc}(\mathscr{K})=\emptyset$, then $\operatorname{Spc}^{cp}(\mathscr{K})=\emptyset$. Explicitly computing $p\operatorname{Spc}$ for sufficiently noncommutative monoidal triangulated categories appears challenging, since it remains unclear when an exact triangle is commutative.